OpenStudy (bdeondrck2):

Find the distance between (-1, 4) and (3, 1)

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -1}}\quad ,&{\color{blue}{ 4}})\quad % (c,d) &({\color{red}{ 3}}\quad ,&{\color{blue}{ 1}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\) you tell us :)

OpenStudy (bdeondrck2):

41? @jdoe0001

OpenStudy (jdoe0001):

what did you get for the radicand?

OpenStudy (mathmale):

BD: Use the distance formula. I'd bet you've seen it before.\[distance~ betw'n ~these~two~points=\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{2})^2}\]

OpenStudy (mathmale):

You are given the two points. What are their x-coordinates? their y-coordinates?

OpenStudy (mathmale):

Substitutte these coordinates into the distance formula. Note that you could Google "distance formula" to find examples of its use.